hamidiso23:

Complete the equation of the line through (1,4) and (2,2)

hamidiso23:

Use exact numbers.

hamidiso23:

solve in y=mx + b format

hamidiso23:

@kjsaif

hamidiso23:

@vocaloid

hamidiso23:

@pixel

Timmyspu:

Well first we have to plug in the points so how can we do that for a y=mx+b equation?

hamidiso23:

-2

hamidiso23:

is the slope right

Timmyspu:

Well for the points it would be (x,y) so with that how would you plug the numbers in?

hamidiso23:

ima use 4 and 1

Timmyspu:

Alright so how would that look in the equation?

hamidiso23:

y - 4 (-2) x - 1

hamidiso23:

is that right

Vocaloid:

slope formula \[m=\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] pick one of your points to be x1,y1, and the other one to be x2,y2. plug in to get the slope. once you have the slope, you can plug in the slope into the point-slope form y - y1 = m(x-x1), once again, plugging in either point to be your x1,y1, then solve for y.

hamidiso23:

no

hamidiso23:

u do 2 -1

Joe348:

lmao thats what i meant

hamidiso23:

-2/1 is the slope

Joe348:

my bad let me rewrite tht

hamidiso23:

@joe348 wrote:
my bad let me rewrite tht
u good

Joe348:



Joe348:

okay there we go

hamidiso23:

I need to find y also

hamidiso23:

answer has to be in y = mx + b format

Joe348:

wouldnt it be y=-2x+1?

AisanDoll: